1/4 of 1/2 of 1/5 of 200:
Answer by RAY100(1637)

Remember "of" means times(*0
.
1/4*1/2*1/5*200/1 = 200/40 =5
.
check
.
.25 * .5 * .2 *200 = 5,,,,ok